The three terms 7x+1, 9x+7, and 3x-4 (not necessarily in order) are 3 consecutive terms of an arithmetic sequence. Find the largest possible value of the common difference of such an arithmetic sequence.
Answer by Edwin McCravy(20060)

Try it all 6 possible ways to order them:

(1) 7x+1, 9x+7, 3x-4 

Common difference = (9x+7)-(7x+1) = (3x-4)-(9x+7)
                        9x+7-7x-1 = 3x-4-9x-7
                             2x+6 = -6x-11
                               8x = -17
                                x = -17/8 = -2.125
  
Common difference = 2x+6 = 2(-2.125)+6 = 1.75

(2) 7x+1, 3x-4, 9x+7

Common difference = (3x-4)-(7x+1) = (9x+7)-(3x-4)
                        3x-4-7x-1 = 9x+7-3x+4
                            -4x-5 = 6x+11
                             -10x = 16
                                x = 16/(-10) = -8/5 = -1.6
  
Common difference = -4x-5 = -4(-1.6)-5 = 1.4

(3) 9x+7, 7x+1, 3x-4

Common difference = (7x+1)-(9x+7) = (3x-4)-(7x+1)
                        7x+1-9x-7 = 3x-4-7x-1
                            -2x-6 = -4x-5
                               2x = 1
                                x = 1/2 = 0.5

Common difference = -2x-6 = -2(0.5)-6 = -7
  

(4) 9x+7, 3x-4, 7x+1

Common difference = (3x-4)-(9x+7) = (7x+1)-(3x-4)
                        3x-4-9x-7 = 7x+1-3x+4
                           -6x-11 = 4x+5
                             -10x = 16
                                x = 16/(-10) = -8/5 = -1.6
                                x = -8/5 = -1.6

Common difference = -6x-11 = -6(-1.6)-11 = -1.4

(5) 3x-4, 7x+1, 9x+7

Common difference = (7x+1)-(3x-4) = (9x+7)-(7x+1)
                        7x+1-3x+4 = 9x+7-7x-1
                             4x+5 = 2x+6
                               2x = 1
                                x = 1/2 = 0.5

Common difference = 4x+5 = 4(0.5)+5 = 7


(6) 3x-4, 9x+7, 7x+1

Common difference = (9x+7)-(3x-4) = (7x+1)-(9x+7)
                        9x+7-3x+4 = 7x+1-9x-7
                            6x+11 = -2x-6
                               8x = -17
                                x = -17/8 = -2.125

Common difference = 6x+11 = 6(-2.125)+11 = -1.75

Case 5 gives the largesr common difference of 7

Edwin